OpenStudy (anonymous):

(3/y)+(6/y)-(1/y)=11

OpenStudy (anonymous):

for y cannot be 0, so just time y at the both side.answer is y=7/11

OpenStudy (anonymous):

sorry 8/11

OpenStudy (king):

(3+6-1)/y=11 8/y=11 y=8/11

OpenStudy (mimi_x3):

\[\large \frac{3}{y}+\frac{6}{y} -\frac{1}{y}=11 \] \[\large \frac{8}{y} =11->11y=8 ->y = \frac{8}{11} \]